# From Language to Action: A Review of Large Language Models as Autonomous Agents and Tool Users

### v1

@@ 9.2 Methodologies and metrics for evaluation
This is often assessed through accuracy or success rates on established benchmarks for coding, like HumanEval, math reasoning datasets such as GSM8K and MATH, and embodied AI tasks in settings like ALFRED and Overcooked-AI[35, 42, 118, 25]. Besides these general benchmarks, evaluations are frequently customized for specific fields, using metrics like discordance and undertriage rates in clinical triage, the emergence of macroeconomic laws in economic simulations, and improvements in material properties for scientific discovery[55, 76, 84].
@@
In addition to these outcomes, a growing number of studies focus on the efficiency and internal reasoning processes of agents. Metrics like token consumption, tool invocation efficiency, and the balance between cost and quality are used to evaluate how resourceful agent solutions are[53, 36]. At the same time, new methods are being developed to investigate the quality of an agent’s reasoning. This includes detailed detection of different types of errors in math problems, assessing Theory of Mind (ToM) accuracy in cooperative games, and evaluating the logical coherence of reasoning sequences using formal graphs[104, 81]. This attention to internal processes is vital for creating more reliable and interpretable agents.
To create a standard evaluation across different tasks, several meta-evaluation frameworks have been proposed. AgentBoard[90] provides a thorough evaluation platform that combines nine unique multi-turn tasks and introduces a detailed progress rate metric for deeper insights beyond simple success rates. Likewise, specialized evaluators like ToolEval[52] have been created to specifically assess tool-use capabilities, measuring metrics like pass rates and solution quality across thousands of real-world APIs. These initiatives indicate a maturation in the field, moving toward more comprehensive and insightful assessment tools that can help develop more capable and aligned agents.
@@ 9.3 Datasets for agent training and grounding
Table 9: Details of the Sixty-Eight publicly available datasets analyzed in this study
Dataset Dataset Type Ref. Dataset Dataset Type Ref.
@@
HumanEval Code Generation [155] MBPP Code Generation [156]
@@
LeetCodeDataset Code Generation [157] CoNaLa Code Generation [158]
@@
RepoBench-P Code Completion [159] Evol-CodeAlpaca Code Instruction [160]
@@
InterCode Interactive Coding [161] GSM8K Math Reasoning [162]
@@
MGSM Math Reasoning [163] Math-Shepherd Math Reasoning [164]
MultiArith Math Reasoning [165] MATH Math Reasoning [166]
@@
TabMWP Math Word Problem [167] MoST Multi-Step Reasoning [168]
STaRK Reasoning [169] BridgeData Reasoning [170]
@@
Aci-bench Agent Interaction [171] HotpotQA Multi-hop QA [172]
@@
MuSiQue Multi-hop QA [173] StrategyQA Multi-step QA [174]
ArxivQA Scientific QA [175] QASPER Scientific QA [176]
@@
NarrativeQA Long-form QA [177] FEVER Fact Verification [178]
@@
QMSum Query Summarization [179] WikiHow Summarization [180]
@@
GovReport Summarization [181] BookSum Long Summarization [182]
@@
Stanford Alpaca Instruction Tuning [183] OpenAssistant Instruction Tuning [184]
@@
OpenOrca Instruction Tuning [185] WildChat Instruction Tuning [186]
@@
IN3 Instruction Tuning [187] FineWeb Pre-training Corpus [188]
@@
RefinedWeb Pre-training Corpus [189] RedPajama Web Text [190]
@@
LLMSecEval Code Security [191] PKU-SafeRLHF Safety [192]
@@
HarmfulQA Safety [193] $H$-RLHF Alignment [194]
@@
BeaverTails Alignment [195] LLM Attacks Adversarial [196]
R2R Embodied AI [197] REVERIE Embodied AI [198]
@@
Touchdown Embodied AI [199] ALFRED Instruction [200]
@@
ALFWorld Instruction [201] TravelPlanner Planning [202]
@@
BlocksWorld Planning [203] HaGRID Gesture Recognition [204]
@@
EgoGesture Gesture Recognition [205] CogScene 3D Scene Understanding [206]
@@
WIDER FACE Face Detection [207] CelebA Face Attributes [208]
@@
CDSL Clinical Data [209] CCLE Genomics [210]
APARENT2 Genomics [211] Saluki Genomics [212]
@@
SKEMPI Protein Binding Affinity [213] PubMed Scientific Abstracts [214]
@@
GAIA General AI Assistant [215] BIG-bench General Evaluation [216]
@@
MMLU General Evaluation [217] CAMEL Multi-Agent Convo. [65]
@@
SOTOPIA-PI Social Simulation [97] ToolBench Tool Use [218]
@@
Mind2Web Web Agent [219] RoboNet Trajectory [220]
@@
A key strategy involves using and adapting existing public benchmarks to build core reasoning and instruction-following skills. Datasets like those for mathematics (GSM8K, MATH), question-answering (HotPotQA, StrategyQA), and coding (HumanEval) are commonly used for fine-tuning models and serve as a baseline for assessing logical and problem-solving abilities[13, 45, 95]. Instruction-tuning datasets, such as Alpaca, are used to improve the overall capabilities of models before tailoring them for specific agent tasks[44]. To promote broader agent learning, there are ongoing efforts to collect large-scale, unified annotations based on various reasoning frameworks across these complex interactive tasks. A notable trend involves using agents to automate the enhancement and diversification of these instructional datasets, creating a constant cycle to improve data quality.
@@
A major development in agent training is creating datasets and simulated environments designed for tool use and interaction in complex digital spaces. ToolBench provides an extensive instruction-tuning dataset with over 16,000 real-world APIs, allowing models to learn to execute advanced instructions and adapt to new tools[93]. For grounding in interactive settings, benchmarks like WebShop and ALFRED serve as established bases for training and testing web navigation and embodied instruction-following, respectively[36, 118]. More intricate environments, such as AndroidArena[113], assess agents on complex tasks that require cooperation between applications, while TIME-ARENA[91] introduces time dynamics and multitasking challenges in simulated household and lab scenarios. AgentBoard[90] combines nine distinct multi-turn, partially-observable tasks from areas like embodied AI, gaming, and web navigation into one evaluation framework.
@@
To ground agents in specialized, high-stakes fields, researchers are curating and generating domain-specific datasets. In the medical sector, this includes datasets derived from clinical triage manuals (TRIAGE AGENT), public EHR data like MIMIC-IV, patient engagement logs, and extensive cancer proteomics resources such as TCGA and CCLE[38, 75, 79]. New methods are being developed to create realistic synthetic Arabic medical dialogues from patient notes, addressing the lack of available data. In science and engineering, agents are trained on tailored datasets drawn from scientific literature for research in organic semiconductors and chemistry or data generated from physics-based simulations for mechanics problems[70, 34]. Similarly, specialized collections are being built for urban planning using public data from cities like San Francisco, for finance using public datasets like MovieLens and Amazon-Beauty to simulate user behavior, and for electronics design automation by generating custom code-instruction pairs[71, 106, 87].
@@
A growing area of research focuses on datasets meant to enhance more nuanced social and collaborative intelligence. The SOTOPIA-$\pi$ dataset supports the interactive learning of social skills through imitation and reinforcement learning based on filtered social interaction data[97]. Other approaches create "societies" of LLM agents to explore emergent collaborative behaviors using established benchmarks like MMLU and MATH[19]. To assess an agent’s ability to recognize implicit human needs, the IN3 benchmark tests proactive clarification skills, while the OEDD corpus presents scenarios where agents must apply different experiential information without being misled by distractions[66, 101]. These varied and increasingly sophisticated datasets are essential for developing agents that not only excel at tasks but are also grounded, reliable, and socially aware.
@@ 10 Discussion
Baseline LLMs. Recent studies reveal a clear hierarchy of foundation models addressing (RQ1) within the landscape of agentic systems. GPT-4 is currently the dominant foundational model, cited in fifty-five studies. Following this, GPT-3.5 has been used in 23 studies, establishing itself as a performance benchmark while introducing cost and access restrictions. GPT variants fulfill three complementary roles: first, as gold-standard benchmarks in ablation studies; second, as collaborators alongside open-source models in multi-agent settings; and third, as primary reasoning modules within complete agent stacks.
@@
In addition to OpenAI models, Anthropic’s Claude 3 series is a prominent proprietary alternative [41, 90, 61, 94]. Google’s Gemini models excel in very-long-context reasoning for document-scale planning tasks. Meanwhile, the open-source ecosystem is rapidly advancing, reducing the performance gap. Models such as Meta’s LLaMA 2/3 [60, 58, 74, 129, 86], Mistral AI’s Mistral series [71, 91, 66, 97, 49, 50, 133], Google’s Gemini [94, 83], Alibaba’s Qwen, DeepSeek, Vicuna, WizardLM, and GLM-4 are increasingly adopted in coding, planning, and dialogue agent tasks. Specialized derivatives like CodeLLaMA and ToolLLaMA demonstrate that adapter-based specialization and task-specific pre-training can confer domain expertise without requiring full retraining. For resource-constrained environments, smaller models, including Phi-2, Phi-3.5 Mini [108], Gemini-2B, and Mistral-7B, offer more efficient options for experimentation and deployment. Lastly, multimodal integration with systems like CLIP, Stable Diffusion XL, and VQ-GAN enables LLM agents to perceive, imagine, and act beyond text, enhancing their capabilities in vision-based and interface automation scenarios.
@@
External tool integration in LLMs. We found that the integration of external tools (RQ2) is one of the principal mechanisms to enable autonomy in LLM agents. Across diverse domains, researchers repeatedly demonstrate that pairing LLMs with rich computational resources, web-search APIs, curated knowledge graphs, code-execution sandboxes, RESTful service libraries [52], and high-fidelity simulators elevates agent capability well beyond what static pre-training LLMs do alone. Real-time data retrieval mechanisms via popular search engines ensure outputs are factually correct when incorporated into RAG systems and supplemented with domain-specific sources. In applied domains, speech recognition tools, such as Speechly, exemplify the potential of real-world interaction. Furthermore, using environments such as AI2-THOR, ALFWorld, and SMAC, connected through platforms like ROS 2 and Gazebo, demonstrates that LLMs can perceive state, reason, and act in dynamic and interactive settings [122, 13, 108, 44, 103]. Collectively, these patterns strongly suggest that external tool access is not just an enhancement feature but an essential foundational mechanism for accessing up-to-date knowledge, enabling perception, action, and reasoning in modern LLM agents.
@@
Frameworks for LLM agents. Our analysis of LLM agent frameworks (RQ3) reveals a clear pattern throughout the literature. ReAct and Reflection dominance in single-agent settings, where streamlined reasoning-action integration and iterative self-improvement are prioritized. ReAct’s integration of thought-action sequences enables dynamic task adaptation, while Reflexion allows agents to refine performance through memory-based learning and feedback incorporation [13]. These frameworks have proven particularly effective for tasks that require adaptability, minimal infrastructure overhead, and self-guided correction.
@@
On the other hand, multi-agent implementations gravitate toward coordination-intensive frameworks such as AutoGen and CAMEL, effective for agent interaction, role differentiation, and collective planning [221]. AutoGen’s conversation programming supports asynchronous and modular communication workflows. CAMEL’s role-conditioned framework enables agents to solve tasks collaboratively through structured, multi-turn dialogues. Both of these capabilities are critical in domains involving multistep reasoning, distributed system design, or collaborative creativity, where autonomous agents operate semi-independently while maintaining coherence with the group objective. LangChain is uniquely positioned in this landscape, supporting chain reasoning steps, memory access, and tool use. It is compatible in single and multi-agent setups [116, 36]. MetaGPT integrates software engineering conventions among less prevalent frameworks, while AIDE and BDI focus on interpretability and logical behaviors to enhance explainability and work in rule-constrained domains.
@@
The trend reflects a departure from static, instruction-bound architectures to adaptive, role-aware, and coordination-driven systems. The widespread adoption of ReAct, Reflexion, LangChain, and AutoGen indicates a convergence on reasoning, memory integration, and collaborative execution as core capacities that underpin the next generation of LLM agents.
@@
Reasoning, planning, and memory. Integration of reasoning, planning and memory mechanisms (RQ4) increasingly shaped the development of LLM-based agents. During analysis, we found distinctions between single-agent and multi-agent systems. Across diverse reasoning approaches, a consistent pattern is: single-agent frameworks emphasize autonomy, safety, and introspective capability, and multiagent frameworks prioritize coordination, belief modeling, and contextual reasoning. For example, TOOLLLM [52], EASYTOOL [53], and Theorem of Thought [49] demonstrate diverse single agent strategies, including search-based decision trees and formal reasoning decomposition.
@@
In contrast, multi-agent frameworks like AGENTVERSE [35] implement collaborative reasoning using structured symbolic modules, and Thread Memory [71] supports contextual retention in multi-turn dialogue and interaction tasks. Theory-of-Mind mechanisms [104] further enhance coordination by modeling belief states in dynamic multi-agent environments. The widely used reasoning methods, CoT, ReAct, and Self-Reflection, appear in both agentic systems. The ubiquity of CoT in tasks such as code generation and robotics [40, 104, 67] underscores its effectiveness in stepwise decomposition. Meanwhile, ReAct [91, 132] and Reflexion [91, 132, 102] integrate feedback-aware reasoning, contributing to real-time adaptability and iterative self-improvement.
@@
Among planning techniques, single agent strategies, such as ReHAC [102] and RAFA [135], emphasize proactive multistep forecasting and heuristic goal decomposition. Multi-agent systems such as DeCoAgent [37] and SMART-LLM [61] address coordination through prompt-based decomposition and coalition formation, respectively. In particular, adaptive and evolution-based planning in Star-Agents [44] and UrbanKGent [86] reflect increasing interest in long-horizon self-organizing workflows.

### v2

@@ 9.2 Methodologies and metrics for evaluation
This is often assessed through accuracy or success rates on established benchmarks for coding, like HumanEval, math reasoning datasets such as GSM8K and MATH, and embodied AI tasks in settings such as ALFRED and Overcooked-AI[34, 41, 114, 25]. In addition to these general benchmarks, evaluations are frequently customized for specific fields, using metrics like discordance and undertriage rates in clinical triage, the emergence of macroeconomic laws in economic simulations, and improvements in material properties for scientific discovery[53, 73, 81].
@@
In addition to these outcomes, a growing number of studies focus on the efficiency and internal reasoning processes of agents. Metrics like token consumption, tool invocation efficiency, and the balance between cost and quality are used to evaluate how resourceful agent solutions are [51, 35]. At the same time, new methods are being developed to investigate the quality of an agent’s reasoning. This includes detailed detection of different types of errors in math problems, assessing ToM accuracy in cooperative games, and evaluating the logical coherence of reasoning sequences using formal graphs[100, 78]. This attention to internal processes is vital for creating more reliable and interpretable agents.
To create a standard evaluation across different tasks, several meta-evaluation frameworks have been proposed. AgentBoard[23] provides a comprehensive evaluation platform that combines nine unique multiturn tasks and introduces a detailed progress rate metric for deeper insight beyond simple success rates. Likewise, specialized evaluators like ToolEval[50] have been created to specifically assess tool-use capabilities, measuring metrics like pass rates and solution quality across thousands of real-world APIs. These initiatives indicate a maturation in the field, moving toward more comprehensive and insightful assessment tools that can help develop more capable and aligned agents.
@@ 9.3 Datasets for agent training and grounding
Table 7: Details of the sixty-eight publicly available datasets analyzed in this study, including their limitations and type.
Dataset Dataset Type Limitations Dataset Dataset Type Limitations
@@
HumanEval [148] Code Generation Limited Task Diversity MBPP [149] Code Generation Limited Concept Coverage
@@
LeetCodeDataset [150] Code Generation Narrow Problem Coverage CoNaLa [151] Code Generation Single Answer Bias
@@
RepoBench-P [152] Code Completion Limited Language Diversity Evol-CodeAlpaca [153] Code Instruction Limited Domain Coverage
@@
InterCode [154] Interactive Coding Limited Language Coverage GSM8K [155] Math Reasoning Limited Problem Variety
@@
MGSM [156] Math Reasoning Limited Problemset Coverage Math-Shepherd [157] Math Reasoning Narrow Problem Scope
MultiArith [158] Math Reasoning Limited Problem Complexity MATH [159] Math Reasoning Lacks Algorithmic Reasoning Problems
@@
TabMWP [160] Math Word Problem Limited Problem Diversity MoST [161] Multi-Step Reasoning Inherent Annotation Noise
STaRK [162] Reasoning Limited Human-Query Variety BridgeData [163] Reasoning Limited Task Diversity
@@
Aci-bench [164] Agent Interaction Synthetic Data Reliance HotpotQA [165] Multi-hop QA Limited Domain Coverage
@@
MuSiQue [166] Multi-hop QA Restricted Question Diversity StrategyQA [167] Multi-step QA Limited Factual Diversity
ArxivQA [168] Scientific QA Limited Disciplinary Diversity QASPER [169] Scientific QA Limited Domain Coverage
@@
NarrativeQA [170] Long-form QA Limited Training Data FEVER [171] Fact Verification Limited Evidence Scope
@@
QMSum [172] Query Summarization Limited Domain Coverage WikiHow [173] Summarization Limited Abstraction
@@
GovReport [174] Summarization Limited Domain Representation BookSum [175] Long Summarization Limited Coverage Of Diverse Genres
@@
Stanford Alpaca [176] Instruction Tuning Limited Contextual Variation. OpenAssistant [177] Instruction Tuning Limited Domain Coverage
@@
OpenOrca [178] Instruction Tuning Limited Domain Coverage WildChat [179] Instruction Tuning Limited Real-World Diversity.
@@
IN3 [63] Instruction Tuning Limited Task Diversity FineWeb [180] Pre-training Corpus Limited Content Diversity
@@
RefinedWeb [181] Pre-training Corpus Contains Residual Noise RedPajama [182] Web Text Benchmark Contamination Risk
@@
LLMSecEval [183] Code Security Limited CWE Coverage PKU-SafeRLHF [184] Safety Limited Scale And Harm Granularity
@@
HarmfulQA [185] Safety Synthetic Prompt Bias HH-RLHF [186] Alignment Limited Generalization Scope
@@
BeaverTails [187] Alignment Narrow Harm Category LLM Attacks [188] Adversarial Limited Attack Diversity
R2R [189] Embodied AI Limited Context Diversity REVERIE [190] Embodied AI Sparse Object Annotations
@@
Touchdown [191] Embodied AI Narrow Urban Context ALFRED [192] Instruction Limited Environment Diversity
@@
ALFWorld [193] Instruction Limited Task Variety TravelPlanner [194] Planning Low Real-World Complexity.
@@
BlocksWorld [195] Planning Lacks Real-World Variability HaGRID [196] Gesture Recognition Lacks Dynamic Gestures
@@
EgoGesture [197] Gesture Recognition Limited Gesture Diversity CogScene [198] 3D Scene Understanding Lacks Scenario Diversity
@@
WIDER FACE [199] Face Detection Demographic Bias CelebA [200] Face Attributes Limited Attribute Diversity
@@
CDSL [201] Clinical Data Lacks Temporal Coverage CCLE [202] Genomics Limited Tissue Representation
APARENT2 [130] Genomics Limited Task Complexity Saluki [203] Genomics Limited Domain Coverage
@@
SKEMPI [204] Protein Binding Affinity Lacks Structural Diversity PubMed [205] Scientific Abstracts Possesses Inconsistent Indexing
@@
GAIA [206] General AI Assistant Lacks Language Diversity BIG-bench [207] General Evaluation Limited Task Diversity
@@
MMLU [208] General Evaluation Data Contains Cultural Bias CAMEL [22] Multi-Agent Convo. Limited Temporal Coverage
@@
SOTOPIA-PI [93] Social Simulation Limited Safety Dimensions ToolBench [209] Tool Use Exhibits Tool Specific Bias
@@
Mind2Web [210] Web Agent Evaluation Bias And Limited Scope RoboNet [211] Trajectory Limited Task Modalities
@@
A key strategy involves using and adapting existing public benchmarks to build core reasoning and instruction-following skills. Datasets like those for mathematics (GSM8K, MATH), question-answering (HotPotQA, StrategyQA), and coding (HumanEval) are commonly used for fine-tuning models and serve as a baseline for assessing logical and problem-solving abilities[13, 43, 91]. Instruction-tuning datasets, such as Alpaca, are used to improve the overall capabilities of models before tailoring them for specific agent tasks[42]. To promote broader agent learning, there are ongoing efforts to collect large-scale, unified annotations based on various reasoning frameworks across these complex interactive tasks. A notable trend involves employing agents to automate the enhancement and diversification of these instructional datasets, creating a constant cycle to improve data quality.
@@
A major development in agent training is creating datasets and simulated environments designed for tool use and interaction in complex digital spaces. ToolBench provides an extensive instruction-tuning dataset with over 16,000 real-world APIs, allowing models to learn to execute advanced instructions and adapt to new tools [89]. For grounding in interactive settings, benchmarks like WebShop and ALFRED serve as established bases for training and testing web navigation and embodied instruction-following, respectively [35, 114]. More intricate environments, such as AndroidArena [109], assess agents on complex tasks that require cooperation between applications, while TIME-ARENA [87] introduces time dynamics and multitasking challenges in simulated household and lab scenarios. AgentBoard [23] combines nine distinct multi-turn, partially observable tasks from areas like embodied AI, gaming, and web navigation into one evaluation framework.
@@
To ground agents in specialized, high-stakes fields, researchers are curating and generating domain-specific datasets. In the medical sector, this includes datasets derived from clinical triage manuals (TRIAGE AGENT), public EHR data such as MIMIC-IV, patient engagement logs, and extensive cancer proteomics resources such as TCGA and CCLE [37, 72, 76]. In science and engineering, agents are trained on tailored datasets drawn from scientific literature for research in organic semiconductors and chemistry or data generated from physics-based simulations for mechanics problems[67, 33]. Similarly, specialized collections are being built for urban planning using public data from cities like San Francisco, for finance using public datasets like MovieLens and Amazon-Beauty to simulate user behavior, and for electronics design automation by generating custom code-instruction pairs[68, 102, 84].
@@
A growing area of research focuses on datasets meant to enhance more nuanced social and collaborative intelligence. The SOTOPIA-$\pi$ dataset supports the interactive learning of social skills through imitation and RL based on filtered social interaction data[93]. Other approaches create "societies" of LLM agents to explore emergent collaborative behaviors using established benchmarks like MMLU and MATH[19]. To assess an agent’s ability to recognize implicit human needs, the IN3 benchmark tests proactive clarification skills, while the OEDD corpus presents scenarios where agents must apply different experiential information without being misled by distractions [63, 97]. These varied and increasingly sophisticated datasets are essential for developing agents that excel in tasks and are grounded, reliable, and socially aware.
Benchmarks and datasets play a crucial role in evaluating LLM-based agents by directly testing key capabilities such as reasoning, planning, tool use, collaboration, and generalization. For reasoning, the GSM8K dataset evaluates mathematical problem-solving. The CORY framework[41] improves policy optimality by reducing distribution collapse, showing strong deductive reasoning. Similarly, HumanEval and MBPP[39] test code generation.
In terms of planning, ALFRED[114] assesses agents few-shot performance . LLM-Planner’s few-shot performance reveals grounding challenges. Similarly, TIME-ARENA[87] examines time-aware multitasking in cooking and lab tasks and highlights LLMs’ difficulties with temporal dependencies.
AGENTVERSE [34] evaluates text understanding and coding, showing emerging cooperative behaviors. SOTOPIA-$\pi$ [93] uses MMLU to achieve social goal completion, matching GPT-4, which emphasizes social intelligence. Tool use is assessed through ToolBench [50], which tests API handling with strong zero-shot generalization. RE-Bench [86] shows that LLMs outperform human experts in time-constrained R&D tasks, indicating effective decision-making.
Multimodal grounding is evaluated through VQAv2 and Text-VQA. CogAgent[57] excels in GUI navigation, which evaluates agents’ visual-language integration. By directly examining the strengths of LLM, these benchmarks also uncover limitations, such as long-term focus issues[40] and social biases[101]. Although datasets in LLM-based autonomous agent research have significantly advanced areas such as reinforcement learning, code generation, healthcare, and urban planning, they still present major limitations that hinder generalization and real-world applicability. Many existing datasets are built within simulated or synthetic environments, including ALFRED[114], Overcooked AI[25], and Sociodojo[88]. These settings often fail to capture the inherent noise, unpredictability, and multimodal complexity of real environments, leading to overfitting and poor performance[109, 58, 48]. Robotics and navigation datasets, for instance Street View[55] and AndroidArena[109] often lack diversity in user behavior, scene variation, and environmental dynamics, resulting in biased models with limited robustness[64, 113]. Standard benchmarks, including HumanEval, MBPP, GSM8K, MATH, and WebShop, remain task-specific with limited cross-domain and multilingual coverage [13, 39, 41, 212].
Existing benchmarks and datasets inadequately measure agentic reasoning by testing performance in simplified, static environments that fail to reflect real-world complexity. They often prioritize subjective, language-level evaluation over assessing an agent’s objective, action-level impact on achieving goals [95]. This reliance on metrics like final success rate offers few insights into the actual reasoning process[23] and neglects crucial challenges like navigating dynamic action spaces[109]. Furthermore, these benchmarks often overlook critical real-world constraints, such as temporal dynamics, which are essential for assessing an agent’s ability to plan and multitask efficiently [87].
@@ 10 Discussion
Baseline LLMs. Recent studies addressing (RQ1) reveal that GPT-4 is currently the dominant foundational model, cited in fifty-five studies, followed by GPT-3.5 in 23 studies, establishing it as a performance benchmark despite cost and access restrictions. GPT variants fulfill three roles: first, as gold-standard ablation references; second, as multi-agent collaborators with open models; and third, as primary reasoning modules within agent stacks.
@@
In addition to OpenAI models, Anthropic’s Claude 3 series is a prominent proprietary alternative [40, 23, 59, 90]. Google’s Gemini excels at very-long-context reasoning for document-scale planning. Meanwhile, the open-source ecosystem is rapidly advancing, reducing the performance gap. Models such as LLaMA-2/3 [58, 56, 71, 124, 83], Mistral [68, 87, 63, 93, 47, 48, 127], Gemini [90, 80], Qwen, DeepSeek, Vicuna, WizardLM, and GLM-4 are increasingly adopted for coding, planning, and dialogue. Specialized derivatives like CodeLLaMA and ToolLLaMA demonstrate, task-specific pre-training can produce domain expertise without complete retraining. For constrained environments, Phi-2, Phi-3.5 Mini [104], Gemini-2B, and Mistral-7B, are more efficient. Lastly, multimodality through CLIP, Stable Diffusion XL, and VQ-GAN enables LLM agents beyond text to vision and interface automation.
@@
External tool integration in LLMs. We found that integrating external tools (RQ2) is a primary driver of LLM autonomy. LLMs paired with rich computational resources, web APIs, knowledge graphs, code execution, RESTful service libraries [50], and high-fidelity simulators elevate agent capability. Real-time data retrieval from domain-specific sources strengthens RAG accuracy while speech recognition tools, such as Speechly, enable real-world interaction. Furthermore, environments such as AI2-THOR, ALFWorld, and SMAC, connected through platforms like ROS 2 and Gazebo, demonstrate that LLMs can perceive state, reason, and act in dynamic and interactive settings [118, 13, 104, 42, 99]. Collectively, these patterns strongly suggest that external tool access is an essential foundational mechanism for accessing up-to-date knowledge, enabling perception, action, and reasoning in modern LLM agents.
@@
Frameworks for LLM agents. Our analysis of LLM agent frameworks (RQ3) reveals single-agent scenario is dominated by ReAct and Reflection, prioritizing streamlined reasoning-action integration and iterative self-improvement. ReAct’s thought-action sequences enable dynamic task adaptation, while Reflexion refines performance through memory-based learning and feedback [13]. These frameworks excel at tasks that require adaptability, minimal infrastructure, and self-guided correction.
@@
Multi-agent implementations favor frameworks such as AutoGen and CAMEL, effective for interaction, role differentiation, and collective planning [213]. AutoGen’s conversation programming supports asynchronous, modular communication workflows. CAMEL’s role-conditioned framework enables agents to solve tasks collaboratively through structured, multi-turn dialogues. These capabilities are critical for multistep reasoning, distributed system design, or collaborative creativity, where agents operate semi-independently while maintaining group objectives. LangChain is compatible in single and multi-agent setups with chain reasoning steps, memory access, and tool use [112, 35]. MetaGPT [21] integrates software engineering conventions among less prevalent frameworks, while AIDE [86] and BDI [99] focus on interpretability and logical behaviors to enhance explainability and work in rule-constrained domains.
@@
The trend is shifting from instruction-bound architectures to adaptive, role-aware, and coordination-driven systems. The widespread use of ReAct, Reflexion, LangChain, and AutoGen indicates a convergence on reasoning, memory integration, and collaborative execution as core capacities.
@@
Reasoning, planning, and memory. Integration of reasoning, planning and memory mechanisms (RQ4) shaped the development of LLM-agents. During analysis, we found single-agent frameworks like TOOLLLM [50], EASYTOOL [51], and ToTh [47] emphasize autonomy, safety, and introspective capability. In contrast, multi-agent frameworks like AGENTVERSE [34], Thread Memory [68], and ToM [100] prioritize coordination, belief modeling, and contextual reasoning. AGENTVERSE implements collaborative reasoning using structured symbolic modules; Thread Memory preserves context in multi-turn dialogue and interaction; ToM enhances coordination by modeling belief states in multi-agent settings. CoT, ReAct, and Self-Reflection appear in both. CoT is effective in code generation and robotics [39, 100], where ReAct [87, 64] and Reflexion [87, 98] integrate feedback-aware reasoning, contributing to real-time adaptability and iterative self-improvement.
@@
Single agent planning strategies, such as ReHAC [98] and RAFA [129], emphasize proactive multistep forecasting and heuristic goal decomposition. Multi-agent systems such as DeCoAgent [36] and SMART-LLM [59] address coordination through prompt-based decomposition and coalition formation, respectively. In particular, adaptive and evolution-based planning in Star-Agents [42] and UrbanKGent [83] reflect increasing interest in long-horizon self-organizing workflows.